If a 30 cm, b 50 cm, Q 5.0 HC, and q-4.0 HC, calculate the magnitude and direction of the electric field (due to these charges) at point P Magnitude: Direction:
Answer #1

electric field due to q ,    E1x = -k*q/b^2 = -9*10^9*4*10^-6/0.5^2 = -144000 N/C


E1y = 0

distance from Q to Point P , r = sqrt(a^2+b^2 )= sqrt(30^2+50^2) = 58.3 cm = 0.583 m

electric field due to Q , E2x = k*Q*b/r^3


E2x = 9*10^9*5*10^-6*0.5/0.583^3


E2x = 113547.3 N/C


E2y = -k*Q*a/r^3


E2y = -9*10^9*5*10^-6*0.3/0.583^3

E2y = -68128.4 N/C


Ex = E1x + E2x = -144000 + 113547.3 = -30452.7 N/C

Ey = E1y + E2y = -68128.4 N/C


magnitude E = sqrt(E^2+Ey^2)


magnitude E = 74624.7 N/C

direction tan^-1(Ey/Ex) = 246 degress counter clockwise with +x axis